# DoPTA: Improving Document Layout Analysis using Patch-Text Alignment

### v1

@@ 4.4 Text Detection
Baselines and Results. We compare against various CNN and ViT backbones and report the precision, recall, and F1 Score at IoU=$0.5$. We do not compare against VGT since it uses the OCR as an input, making the task redundant. StructTextv2 is omitted due to non-availability of code/weights. DoPTA-HR outperforms the previous best result from DiT-L, despite pre-training for only $1/2$ the pre-training steps, and having less than $1/3^{rd}$ the parameters.
@@ 5 Ablation Study
In this section, we study the effect of the individual components of DoPTA. It is crucial to study the performance of our proposed patch-text contrastive loss. The reconstruction loss is also an important component, bringing visual information to the features where textual features are not available. To understand the contribution of each loss objective independently, we evaluate our method on an array of different masking ratios, as well as in the absence of the reconstruction loss. We also compare with the CLIP model, to quantify the contribution of the architecture, in the absence of our loss components. All experiments in the ablation study were carried out at $224$ resolution while pre-training for 10 epochs on the IIT-CDIP Dataset. The batch size, learning rate, data augmentation, and other hyperparameters were kept the same as the original setup. We evaluate the performance by benchmarking on document image classification and document layout analysis. The setup and fine-tuning parameters for each downstream evaluation are unchanged from Sec. 4.
Results. The results of the ablation study are summarized in Table 4. It is clear that both loss components signifcantly improve performance over the Row 1 baseline. Further, as seen in Row 2, the model trained only with the proposed patch-text contrastive loss retains strong performance. In particular, this variant outperforms DiT-Base on RVL-CDIP, while almost matching performance on PubLayNet, despite the comparatively short pre-training schedule adopted in this study (10 epochs in ours as compared to 30 epochs in DiT). This result highlights the efficacy of the patch-text contrastive loss in learning effective visual representations for document layout analysis. The results see a clear improvement when the reconstruction loss is included. We see an improvement of $0.3-0.6$ performance points in all benchmark scores as the masking ratio is increased. The combination of patch-text alignment and reconstruction objectives enables the model to learn strong visual representations that generalize across various task settings.
@@ 6 Conclusion and Future Work
In this work, we extend fine-grained image-text alignment to document images via a novel patch-text alignment objective. This work shows the efficacy of leveraging the textual information in document images to solve visual tasks, which is still an underexplored direction. We combine this novel objective with a masked reconstruction loss to build DoPTA, a strong document encoder model that achieves state-of-the-art results across document image classification, layout analysis, and text detection tasks, consistently outperforming baselines that use larger models, extra information (OCR) as input, and longer pre-training schedules. We hope that this work motivates further research into methods that can leverage text in document images for visual understanding.
There still remain many open avenues for exploration. A few of these include - Extending DoPTA to newer architectures like SwinTransformer which are better suited for document images with small objects and details, alternative strategies for text masking, and leveraging synthetic data generation techniques to increase the size and diversity of the training dataset. We aim to explore these directions as our future work.
@@
The appendix is structured as follows - In Appendix A we present and discuss results of pre-training DoPTA on the PixParse dataset. Appendix B contains additional qualitative examples of the effect of our various pre-training objectives. In Appendix C we qualitatively analyse the relative lower performance of DoPTA-HR on certain categories in D4LA. In Appendix D we present the inference time comparison of our detection model.
@@ Appendix A Results on PixParse
In Table 6, we report the results of pre-training DoPTA on the PixParse dataset, and compare it to the variant trained on CDIP. All hyperparameters for DoPTA and DoPTA-HR are identical to the original pre-training setting outlined in Sec. 4. However, we only train for $80$K steps as we find this sufficient to notice significant differences between pre-training on PixParse and CDIP. We notice a consistent trend, wherein DoPTA pre-trained on PixParse achieves consistently lower scores than the CDIP variant across all benchmarks. Despite the high quality OCR data, this may be caused due to to major reasons - i) The low number of samples in PixParse, leading to a larger degree of overfitting on the pre-training dataset, and ii) The distribution of downstream benchmarks like RVL-CDIP and FUNSD more closely matching that of the pre-training data in CDIP, as these are both datasets comprising of scanned documents, which may prove to be slightly OOD when pre-training on PixParse.
@@
 DoPTA CDIP 92.88 67.71 91.96
 DoPTA PixParse 92.28 67.33 91.96
 DoPTA-HR CDIP 93.78 69.69 94.31
 DoPTA-HR PixParse 93.47 68.77 94.19
@@
MODEL Inference time (s)
VGT 34.27
 DoPTA-HR 2.06
Table 7: Comparison of inference time per sample (in sec) of our DoPTA  vs VGT on performing objection detection oh D4LA using the Cascade RCNN framework. Note that this time excludes the additional inference time required by VGT to perform OCR extraction.
@@ Appendix B Additional Qualitative Examples
In Fig. 4, we demonstrate additional qualitative examples of the effect of the patch-text alignment objective on the DoPTA encoder. The results demonstrate the ability of the pretrained DoPTA -HR encoder to isolate individual words in a document image, which was the goal of the patch-text alignment objective. This ability translates to better performance on downstream benchmarks, as demonstrated in Sec. 4.
@@
Figure 4: Heatmap visualisation of the normalised dot product similarity of image patch embeddings and text embeddings taken from the DoPTA-HR model, demonstrating its ability to find individual words in document images. Despite some noise, there is a clear spike in dot product similarity at the appropriate text region. The target text word for each image is mentioned.
Appendix C Analysis of Failure Cases
In this section, we analyze the lower performance of DoPTA-HR on certain classes in the D4LA dataset. In particular, we observe lower performance on the RegionList category. We found that this occurs due to a common error made by DoPTA-HR, as demonstrated in Fig. 5, where the model incorrectly marks RegionList as RegionKV. This is most likely due to the high visual similarity between the two classes, and the ground truth labels often seem to be ambiguous. Another area of low performance was the Equation category, where DoPTA-HR ($32.26$ mAP) yields far lower performance than VGT ($49.0$ mAP). However, this category has an extremely low occurence in the dataset, which may explain the low performance. We did not observe any other consistent trend which may explain said low performance.
Appendix D Timing Analysis
In Table 7 we analyze the average running time for a single image across different models. VGT has a significantly longer running time, owing to the larger model size, despite having pre-extracted OCR. Despite the faster running time, DoPTA-HR outperforms VGT consistently on multiple benchmarks, as demonstrated in Sec. 4.
Figure 5: Failure case of DoPTA-HR on layout analysis on D4LA benchmark. Left is DoPTA-HR. Right is VGT. DoPTA-HR incorrectly marks the central region as RegionKV, which was found to be a common error mode.

### v2

@@ 4.4 Text Detection
Baselines and Results. We compare against various CNN and ViT backbones and report the precision, recall, and F1 Score at IoU=$0.5$. We do not compare against VGT since it uses the OCR as an input, making the task redundant. StructTextv2 is omitted due to non-availability of code/weights. DoPTA outperforms the previous best result from DiT-L setting new SOTA, despite pre-training for only $1/2$ the pre-training steps, and having less than $1/3^{rd}$ the parameters.
4.5 Inference Time Analysis
In this section, we analyze the performance benefits of our OCR-free inference setting compared to baselines on layout detection models on the D4LA test set. We observe that OCR parsing with EasyOCR[21] takes an average of 1.02 seconds per image. Since VGT relies on OCR during inference, it operates at 0.81 FPS while our  DoPTA  achieves 9.56 FPS which is —12× faster than VGT. Not only does  DoPTA  achieve SOTA performance, but it also has significantly improved inference speed. A visual comparison is provided in Figure 6. All the methods were carefully tested on the same A100 gpus.
@@ 5 Ablation Study
In this section, we study the effect of the individual components of DoPTA. It is crucial to study the performance of our proposed patch-text contrastive loss. The reconstruction loss is also an important component, bringing visual information to the features where textual features are not available. To understand the contribution of each loss objective independently, we evaluate our method on an array of different masking ratios, as well as in the absence of the reconstruction loss and patch-text contrastive loss. We also compare with the CLIP model, to quantify the contribution of the architecture, in the absence of our loss components. All experiments in the ablation study were carried out while pre-training for 10 epochs on the IIT-CDIP Dataset. The batch size, learning rate, data augmentation, and other hyperparameters were kept the same as the original setup. We evaluate the performance by benchmarking on document image classification and document layout analysis. The setup and fine-tuning parameters for each downstream evaluation are unchanged from Sec. 4.
Results. The results of individual loss components are summarized in Table 5 and the effect of masking ratio in Table  6. It is clear that both loss components signifcantly improve performance over the Row 1 baseline. Further, as seen in Row 3, the model trained only with the proposed patch-text contrastive loss retains strong performance. In particular, this variant outperforms reconstruction only training on RVL-CDIP and D4444LA, while matching performance on PubLayNet. This result highlights the efficacy of the patch-text contrastive loss in learning effective visual representations for document layout analysis. The results see a clear improvement when both the losses are included. In Table 6 we see an improvement of $0.3-0.6$ performance points in all benchmark scores as the masking ratio is increased. The combination of patch-text alignment and reconstruction objectives enables the model to learn strong visual representations that generalize across various task settings.
@@ 6 Conclusion and Future Work
In this work, we extend fine-grained image-text alignment to document images via a novel patch-text alignment objective. Our work shows the efficacy of leveraging the textual information in document images to solve visual tasks, which is still an underexplored direction. We combine this novel objective with a masked reconstruction loss to build DoPTA, a strong document encoder model that achieves state-of-the-art results across document image classification, layout analysis, and text detection tasks, consistently outperforming baselines that use larger models, extra information (OCR) as input, and longer pre-training schedules. We hope that this work motivates further research into methods that can leverage text in document images for visual understanding.
Our work opens several new avenues for further exploration. We aim at extending  DoPTA to newer architectures such as SwinTransformer, which could provide better results for document images with small objects and details, exploring alternative strategies for text masking, and leveraging synthetic data generation techniques to increase the size and diversity of the training dataset. We aim to explore these directions as part of our future work.
@@
The appendix is structured as follows - In Appendix A we present and discuss results of pre-training DoPTA on the PixParse dataset. Appendix B contains additional qualitative examples of the effect of our various pre-training objectives. In Appendix C we qualitatively analyse the relative lower performance of DoPTA on certain categories in D4LA.
@@ Appendix A Results on PixParse
In Table 8, we report the results of pre-training DoPTA on the PixParse dataset, and compare it to the variant trained on CDIP. All hyperparameters for DoPTA  are identical to the original pre-training setting outlined in Sec. 4. However, we only train for $80$K steps as we find this sufficient to notice significant differences between pre-training on PixParse and CDIP. We notice a consistent trend, wherein DoPTA pre-trained on PixParse achieves consistently lower scores than the CDIP variant across all benchmarks. Despite the high quality OCR data, this may be caused due to to major reasons - i) The low number of samples in PixParse, leading to a larger degree of overfitting on the pre-training dataset, and ii) The distribution of downstream benchmarks like RVL-CDIP and FUNSD more closely matching that of the pre-training data in CDIP, as these are both datasets comprising of scanned documents, which may prove to be slightly OOD when pre-training on PixParse.
@@
 DoPTA CDIP 93.78 69.69 94.31
 DoPTA PixParse 93.47 68.77 94.19
@@ Appendix B Additional Qualitative Examples
In Fig. 9, we demonstrate additional qualitative examples of the effect of the patch-text alignment objective on the DoPTA encoder. The results demonstrate the ability of the pretrained DoPTA  encoder to isolate individual words in a document image, which was the goal of the patch-text alignment objective. This ability translates to better performance on downstream benchmarks, as demonstrated in Sec. 4.
Appendix C Analysis of Failure Cases
In this section, we analyze the lower performance of DoPTA on certain classes in the D4LA dataset. In particular, we observe lower performance on the RegionList category. We found that this occurs due to a common error made by DoPTA , as demonstrated in Fig. 5, where the model incorrectly marks RegionList as RegionKV. This is most likely due to the high visual similarity between the two classes, and the ground truth labels often seem to be ambiguous. Another area of low performance was the Equation category, where DoPTA  ($32.26$ mAP) yields far lower performance than VGT ($49.0$ mAP). We identify that this category has an extremely low occurence in the dataset (only 2/3 samples in total), which may explain the low performance. We did not observe any other consistent trend which may explain said low performance.  DoPTA  however demonstrates interesting performance on equation class as illustrated in Fig.8 and 7 where we observe  DoPTA predicting equation entities missed by VGT but with incorrect boundaries leading to poor performance.  DoPTA also predicts chemical equation which was missed by the VGT. We identified that the very high class imbalance in the total dataset led to high variation in the final mAP performance.
Figure 5: Failure case of DoPTAon layout analysis on D4LA benchmark. Left is DoPTA. Right is VGT. DoPTAincorrectly marks the central region as RegionKV, which was found to be a common error mode.
Figure 6: Plot explaining FPS and Publaynet accuracy of various models. Model(OCR) denotes the FPS when OCR parsing is taken into account for computing inference time.
Figure 7: Prediction from  DoPTAon layout analysis on Equation class. Left is DoPTA. Right is VGT. DoPTAidentifies equation objects that were not identified by VGT but also encloses extraneous regions leading to poor performance.
Figure 8: Prediction from  DoPTAon layout analysis on Equation class. Left is DoPTA. Right is VGT. DoPTAidentifies a chemical equation objects that were not identified by VGT.
@@ Appendix C Analysis of Failure Cases
Figure 9: Heatmap visualisation of the normalised dot product similarity of image patch embeddings and text embeddings taken from the DoPTA model, demonstrating its ability to find individual words in document images. Despite some noise, there is a clear spike in dot product similarity at the appropriate text region. The target text word for each image is mentioned.